Ronda Manufacturing Corporation uses a standard cost system with machine-hours as the activity base for overhead. Last year, Ronda incurred $840,000 of fixed manufacturing overhead and generated a $42,000 favorable fixed manufacturing overhead budget variance. The following data relate to last year’s operations:

Denominator activity level in machine-hours 21,000
Standard machine-hours allowed for actual output 20,000
Actual number of machine-hours incurred 22,050

What amount of total fixed manufacturing overhead cost did Ronda apply to production last year?

Multiple Choice:

A) $837,900

B) $840,000

C) $926,100

D) $972,405
Answer #1

Correct answer will b C) $926100

Working for your easy understanding

Total Budgeted Manufacturing Overhead= Actual Overhead + FMOH Variance
=840000+42000=882000
Activity Level in Machine Hour= 21000
Budgete MOH per Hour= 882000/21000=$42
Actual No. of Machine Hour Incurred= 22050
Total Applied Manufactuing Overhead= Actual Machine Hour X Budgeted MOH per Hour
=22050 Machine Hour X $42= $926100